OpenStudy (anonymous):

The measures of all angles in a regular octagon are (4x + 12)°. Find the value of x. Show equations and all work that leads to your answer

OpenStudy (campbell_st):

well the angle sum of a polygon is Angle Sum = 180(n -2) where n = the number of sides in the polygon A regular polygon has all sides and angles equal in their measurements. so to find the value of x you need \[4x + 12 = \frac{anglesum}{8}\] hope it helps